First question:
Right now, Alan's mother is 4 times older than he is. In 4 years, she will be 3 times older than he is. How old are Alan and his mother right now?
Second question:
Kim is using a list of his friends' ages. He adds two of his friends' ages and gets 17. He multiplies the ages and gets 72. How old are his two friends?

I would appreciate your help. Thanks.
-- Mary

Answer by AnlytcPhil(1806)

First question:
Right now, Alan's mother is 4 times older than he is. In
4 years, she will be 3 times older than he is. How old are
Alan and his mother right now?
`
Let Alan's age be X
`
>>...Alan's mother is 4 times older than he is...<<
`
Mother's age = 4X
`
>>...In 4 years,...<<
`
` (Alan's age in 4 years) = X+4
(Mother's age in 4 years) = 4X+4
`
>>...In 4 years, she will be 3 times older than he is...<<
`
(Mother's age in 4 years) = 3 times (Allen's age in 4 years)
`
` ` ` ` ` ` ` ` ` ` `4X+4 = 3(X+4)
`
Can you solve that? Answer: X = 8, Mother's age = 4×8 = 32.
`
Check:
>>...Right now, Alan's mother is 4 times older than he is...<<
`
That's true because he is 8 and she is 4 times 8 or 32.
`
>>...In 4 years, she will be 3 times older than he is...<<
`
That's true because he will be 12 and she will be 3 times
12 or 36.
`
-----------------------------------
`
Second question: Kim is using a list of his friends' ages.
He adds two of his friends' ages and gets 17. He multiplies
the ages and gets 72. How old are his two friends?
`
Let x be one friend's age and y be the other friend's age:
`
>>...He adds two of his friends' ages and gets 17...<<
`
Translation:` x + y = 17
`
>>...He multiplies the ages and gets 72...<<
`
Translation:` ` `xy = 72
`
You have this system to solve
`
` ` ` ` ` ` ` x + y = 17
` ` ` ` ` ` ` ` `xy = 72
`
Solve the first equation for y
`
` ` ` ` ` ` ` ` ` y = 17-x
`
Substitute 17-x for y in the second equation:
`
` ` ` ` ` ` ` ` ` xy = 72
` ` ` ` ` ` `x(17-x) = 72
` ` ` ` ` ` 17x - x² = 72
` ` ` -x² + 17x - 72 = 0
` ` ` `x² - 17x + 72 = 0
` ` ` (x - 9)(x - 8) = 0
` ` ` x - 9 = 0 or x = 9
` ` ` x - 8 = 0 or x = 8
`
`If x = 8, then y = 17-8 or 9
`If x = 9, then y = 17-9 or 8
`
`Either way, one of them is 8 and the other 9.
`
`Edwin